OpenStudy (anonymous):

Find the LCD. 2/x+1 - 3/x-1 =9 9(x + 1)(x - 1) x2 + 1 (x + 1) + (x - 1) (x + 1)(x - 1)

TheSmartOne (thesmartone):

\[\frac{ 2 }{ x+1}-\frac{ 3 }{ x-1 }=9\]

TheSmartOne (thesmartone):

We need the same denominator so the denominator should be (x+1)(x-1)

TheSmartOne (thesmartone):

Which is D.

OpenStudy (anonymous):

Thank you so much!!

TheSmartOne (thesmartone):

You are welcome!